OpenStudy (anonymous):

What is the period of the function p(x) = 2cos(0.5x)?

OpenStudy (anonymous):

\[0.5 x=2 \pi,\frac{ 5 }{ 10 }x=2 \pi,x=2 ~\pi*\frac{ 10 }{ 5 }=4 \pi\] period=?